# Language Models Can Predict Their Own Behavior

### v1

@@ 6 Analysis and Ablations
Figure 5: Left: Density plot of the correlation coefficient between the CoT model accuracy and the consistency of the probe estimate. The coefficient is often positive, indicating that instances where the probe inaccurately estimates CoT model are also instances where the model is more likely to be incorrect. Right: On CosmoQA, probes have near-random consistency with initial tokens. There is a sudden performance boost towards the final token. Confidence remains steady, suggesting poor calibration in early token probes.
@@
Why does accuracy improve when using probes? Seeking to explain the surprising increase in accuracy when accelerating CoT models (Figure 3), we plot (Figure 5) the correlation between the CoT models accuracy and the probes estimation consistency. It is generally a positive correlation, suggesting that on the instances where the probe estimates incorrectly, the CoT model is also more likely to output an incorrect answer. This minimizes the harm of an inconsistent estimation as sometimes making an inaccurate estimate is actively beneficial for overall accuracy.
@@
How early in the computation does the model show signs of the final behavior? We have shown that the final input token often contains sufficient information to predict output behavior, but at which token does this information start becoming clear and accessible? To investigate this, we collect the activations of the mid-layer from every input token (after the FewShot example tokens) and train a linear classifier to map any of these internal activations to the desired output behavior.
Probing (Figure 5) using the embeddings of tokens in the first quarter of the question tokens leads to near-random performance. While consistency increases steadily as we use later tokens, a considerable jump is noticed when shifting to the final tokens. Surprisingly, confidence declines marginally as we use later tokens for probing. This suggests that using the earliest tokens for probing could result in inconsistent and poorly calibrated models, making it a poor design decision for this method.
How does output length affect estimation accuracy? We measure the correlation between the estimation consistency of the probe and the token count of both the input (-0.41%) and output (-6.1%). While input length seems to have little effect, a higher output length is correlated with worse accuracy. This suggests probes struggle more with inputs that evoke longer outputs.
Figure 6: Probes achieve high accuracy with as few as 1000 datapoints in most cases
How much data does a probe need? We ablate the amount of data used to train the probe. We find (Figure 6) that often, fewer than 1000 datapoints is sufficient for the probe to achieve high estimation accuracy, suggesting that the method is particularly data efficient.
Figure 7: Layer ablations show that early layers perform poorly, while the decision between mid and late layers is dataset specific
@@
Which layer should we use to probe? We explore (Figure 7) the specific layer being probed affected the performance of the probe. The most general finding is that the early layers offer poor estimation consistency. However, whether mid or late layers perform better is task and dataset-specific, and cannot be described generally. For more details see Appendix B.
Figure 8: Increasing the confidence threshold has no effect until a point, after which conformal consistency increases while the coverage tends to zero
How does varying $\alpha$ change performance? Increasing the confidence required from the probes has predictable effects (Figure 8). Increasing the confidence threshold has no effect until a point, after which conformal consistency increases while the coverage tends to zero.
Figure 9: Probe accuracy increases with model size, suggesting the method may scale favorably
@@
How does model scale affect performance? We varied the size of the LM used (Llama3 3B, 8B and 70B) and measure probe consistency across several layers. Encouragingly (Figure 9), the performance of the larger models are consistently better, suggesting that the information encoded in the internal activations is easier for the probe to ‘read’ when the model is more powerful. This bodes well for the methods’ ability to scale with models of increasing size and capability.
@@
What are the limits of the probes? Using the MCQ task, we tried to estimate whether the LM would output an incorrect answer. The probe accuracies were consistently poor and near random. This shows that, intuitively, the probes have limitations in the behaviors they can estimate. We hypothesize that since probes are simple linear classifiers, they can only detect patterns and use ‘knowledge’ that is well encoded in the internal activations being probed. This suggests that probes struggle with output properties that cannot be identified by the LM itself (without external knowledge).
@@ 7 Related Works
This work takes inspiration from recent methods that probe the hidden states of LMs to observe interpretable patterns [3, 35, 50, 23, 2] identify false statements [7, 39, 40, 69] and hallucinations [12, 57, 29]. Our work also has connections to literature on early exiting during the forward pass of NN models, with works often using signals from the hidden states to prematurely exit with a prediction on the next token [65, 70, 66, 54, 27]. Notably,  Schuster et al. [54] also uses a conformal prediction framework to give a provable error bound on the next token approximation.
We are the first to show that the internal states can predict a range of behaviors, uninferable from the next token alone, before any of the output tokens are generated. Additionally, we are the first to show that conformal prediction can be used to create early warning systems for a wide range of behaviors like question abstention to format following errors. Our work advances research on understanding the nature of the information contained in the hidden states of LMs [50, 5, 47, 41, 61]. Specifically, we show that the information contained in the hidden states is relevant not just to the next token, but to behaviors that manifest several tokens later during the LMs generation.
@@ 8 Conclusion
We show that a Language Model’s internal representation of input tokens alone contains vital information on the behavior of the LM over the entire output sequence. These signals can be used to create probes that serve as precise early warning or exit systems for LMs on a wide range of behaviors. On 27 text classification datasets across 5 different tasks, the method can accelerate Chain-of-Thought prompting by 65% with little accuracy loss. The method can be used to create high-precision early warning systems that identify if an LM will refuse to answer a question, fall victim to jailbreaking, fail to follow output format specifications or give low-confidence responses. We show that the probes generalize to out-of-distribution test sets, and scale favorably to larger LMs. Finally, we explore the limitations of the method, showing that the behavior of longer output sequences is harder to estimate and that tasks that require knowledge external to the model are particularly challenging. With the rising popularity of inference-time scaling methods, we hope our work can help ameliorate the growing computational cost of running LMs and provide more insight into the nature of the information contained in the hidden states of LMs.
@@ A.1 Confirming Robustness to Choice of Model
We conduct experiments on two other Language Models, to ensure that our results hold on Language Model families outside of the Llama3 series. We use Mistral-7B-Instruct-v0.3 from MistralAI [28] and DeepSeek-R1-Distill-Qwen-14B from DeepSeek [20]. The results show that the phenomenon shown in the main text can be detected robustly across various model families.
@@ A.2 Data Leakage
Several datasets used in this study have been released before the creation of Llama3, and as such might have been pre-exposed to the model during training. We detail our reasons as to why data leakage is not a concern for the conclusions that we draw from this study:
  • There are datasets in our study that have not been leaked. For example, the MMLU test set is used as an official test benchmark by LLama3[16] and WildJailbreak was released to the public only after the initial release of Llama3 models [30]. The results are strong on such datasets as well, showing that data leakage is not behind the performance of the probes.
  • Several tasks require the probes to identify behavior that fundamentally does not exist in text form on the internet. The bullets, JSON and Confidence tasks require the probes to determine how the model will behave, regardless of whether the model has seen the input text or not. For example, whether or not the model was pretrained on MSMarco questions is immaterial for predicting whether or not it will output answers in a specific JSON format.
@@ Appendix B Extended Analysis
B.1 Layer Ablation
@@ C.1.5 Toxicity Detection:
We use two datasets provided by the same organization, JigsawToxicity  and a JigsawUnintendedBiasToxicity . Both datasets have scores for toxicity from a chatroom setting.
@@ C.2.1 Text Classification Tasks:
MCQ: For the MCQ tasks, we select two options by randomly sampling an incorrect answer from the provided options along with the correct option. This is because the MCQ datasets vary in terms of number of MCQ options, and for the out-of-distribution experiment we require all the datasets to have the same number of answer classes. The CoT model is asked to identify the correct answer option.
@@ C.2.2 Other Tasks:
LM Abstention: For the unanswerable question datasets of SelfAware and KnownUnknown, the model is first instructed to give a reasoning chain about the question, and answer only if the question is answerable (abstaining otherwise). With WildJailbreak, the input need not be a question, but any request. Hence, we instruct the model to comply with the request if it is not malicious, and abstain otherwise. In all cases, the probe is trained to detect whether or not the LM will abstain.
@@ C.3 Prompts:
Each task and dataset has a different set of prompts, we have provided a link to the code https://github.com/DhananjayAshok/LMBehaviorEstimation. All prompts can be seen in the file data.py
@@ C.4 Hardware:
All of our experiments were run on a compute cluster with 8 NVIDIA A40 GPUs (approx 46068 MiB of memory) on CUDA version 12.6. The CPU on the cluster is an AMD EPYC 7502 32-Core Processor. Most experiments could be conducted with less than 16GB of RAM.

### v2

@@ 6 Analysis and Ablations
Why does accuracy improve when using probes? Seeking to explain the surprising increase in accuracy when accelerating CoT models (Figure 4), we plot (Figure 5, left) the correlation between the CoT model’s accuracy and the probe’s estimation consistency. It is generally positive, which suggests that incorrect probe estimation correlates with incorrect CoT generation. This minimizes the harm of an inconsistent estimation of the CoT decision, as the cases where estimates are likely to be inconsistent coincide with cases where a consistent estimate does not benefit accuracy.
Figure 5: Left: Density plot of the correlation coefficient between the CoT LM accuracy and the consistency of the probe. The coefficient is often positive, indicating that instances where the probe inaccurately estimates the LM decision are also instances where the LM is more likely to be incorrect. Right: On CosmoQA, probes that use the initial tokens of the input prompt have near-random consistency, with performance increasing considerably as we use later tokens in the prompt.
@@
How early in the computation does the model show signs of the final behavior? We have shown that the final input token often contains sufficient information to predict output behavior, but at which token does this information start becoming clear and accessible? To investigate this, we collect activations from every input token (after the FewShot example tokens) of the CSQA dataset and train a linear classifier to map these internal activations to the eventual CoT prediction. Probing (Figure 5, right) using the embeddings of tokens in the first quarter of the question leads to near-random performance, showing that this information is not readily accessible at initial tokens. Consistency increases steadily as we use later tokens, with a sharp rise around the final tokens.
Figure 6: Left: Ablations on MMLU. Within 500 training datapoints, probes ($\alpha=0.9$) often approach the consistency and coverage they will attain when trained on the entire dataset. Layer ablations ($\alpha=0.9$) show that early layers perform poorly. Increasing the confidence threshold $\alpha$ has no effect until a point, after which conformal consistency increases while the coverage tends to zero. Right: Estimation consistency increases with LM size, suggesting the method may scale favorably
Ablations on probe parameters: We ablate the parameters involved in the design of the conformal probing system: number of training datapoints, probing layer and the conformal confidence threshold $\alpha$. We show results (Figure 6, left) for the MMLU dataset, however, all trends discussed hold for the majority of the datasets we explored, with more examples in Appendix B.
@@
We ablate the amount of data used to train the probe and find (Figure 6, top left) that probes can approach the consistency and coverage they will achieve over the entire dataset with only 500 datapoints, suggesting that the method is particularly data efficient. We explore (Figure 6, middle left) whether the specific layer being probed affects the performance of the probe. Our most general finding is that early layers offer poor estimation consistency. However, whether mid or late layers perform better is task and dataset-specific, and cannot be described generally. For more details, see Appendix B. Increasing the confidence required from the probes has predictable effects (Figure 6, bottom left). Increasing the confidence threshold has no effect until a point, after which conformal consistency increases while the coverage tends to zero.
@@
How does model scale affect performance? We varied the size of the LM used (Llama3 3B, 8B and 70B) and measured probe consistency across several layers. Encouragingly (Figure 6, right), the performance of the larger models is consistently better, suggesting that the information encoded in the internal activations is easier for the probe to ‘read’ when the model is more powerful. This bodes well for the methods’ ability to scale with models of increasing size and capability.
@@
What are some limitations of the probes? Other experiments we conducted suggest that probes are limited in the behaviors they can detect. Using the MCQ task, we tried to estimate whether the LM would output an incorrect answer. The probe accuracies were consistently near random. We hypothesize that since probes are simple linear classifiers, they can only detect patterns and use ‘knowledge’ that is well encoded in the LM activations. This suggests that probes struggle with output properties that cannot be identified by the LM itself (without external knowledge). We also observed that the correlation between the estimation consistency of the probe and the token count of the output is negative (-6.1%). This suggests probes struggle more with inputs that evoke longer outputs.
@@ 7 Related Works
This work takes inspiration from recent methods that probe the hidden states of LMs to observe interpretable patterns [4, 51, 71, 36, 3] identify false statements [10, 56, 57, 97] and hallucinations [20, 84, 44]. These works use the internal states of every token, even the outputs, to produce signals for these phenomena. In contrast, our work shows that internal states can predict behaviors  before any output tokens are generated, suggesting that input token embeddings contain rich information on future LM behavior. Furthermore, while conformal prediction has been used with LMs to guarantee the quality of generated text [73], robotic trajectories [89, 74] and text classification [52, 17], we are the first to use it in conjunction with hidden state probes to create precise early warning systems.
Our work also has connections to the literature on early exiting during the forward pass of NN models, with works often using signals from the hidden states to prematurely exit with a prediction on the future outputs [93, 99, 94, 80, 42, 70]. While these works focus on the tokens predicted, we show a more general result: that internal states can predict the future behavior of LMs. These include behaviors such as whether or not the LM will output a high perplexity answer, or mistakenly comply with a malicious request, which cannot be inferred just by observing the output token sequence.
While recent work has begun exploring whether LMs exhibit ‘introspective’ properties in the black box setting [13], or whether LMs can predict ‘global attributes’ of their response [72, 25], we are the first to show that when internal probes are deployed under conformal prediction they can be used to create early warning systems for a wide range of behaviors like question abstention to format following errors. Our work advances research on understanding the nature of the information contained in the hidden states of LMs [71, 6, 66, 58, 87, 62]. Specifically, we show that the information contained in the hidden states is relevant not just to the next token, but to behaviors that manifest several tokens later during the LMs generation.
@@ 8 Conclusion
We show that a language model’s hidden representation of input tokens alone contains vital information on the behavior of the LM over the entire output sequence. We train linear probes to read this information and use it in a conformal prediction framework to create precise early warning or exit systems for a wide range of LM behaviors, including degenerate behaviors, safety alignment failures and more. The conformal probes can preemptively identify, before a single token is generated, instances where an LM will fail to abstain from answering an unanswerable question, fall victim to jailbreaking, fail to follow output format specifications or give low-confidence responses. On 27 text classification datasets across 5 different tasks, the method can accelerate Chain-of-Thought prompting by 65% with little accuracy loss. We show that the probes generalize to out-of-distribution test sets and scale favorably to larger LMs. Finally, we explore the limitations of the method, showing that the behavior of longer output sequences is harder to estimate and that tasks that require knowledge external to the model are particularly challenging. With the rising popularity of inference-time scaling methods, we hope our work can help ameliorate the growing computational cost of running LMs and provide more insight into the nature of the information contained in their hidden states.
Acknowledgments
We thank Robin Jia for his exceptional course on The Science of Large Language Models (accessible at https://robinjia.github.io/classes/fall2024-csci699.html) and his advice on the early stages of this project. We also acknowledge support from Open Philanthropy.
@@ A.1 Confirming Robustness to Choice of Model
We conduct experiments on two other language models to ensure that our results hold for language model families outside of the Llama3 series. We use Mistral-7B-Instruct-v0.3 from MistralAI [43] and DeepSeek-R1-Distill-Qwen-14B from DeepSeek [32]. The results show that the phenomenon shown in the main text can be detected robustly across various model families.
@@ A.2 Data Leakage
Several datasets used in this study have been released before the creation of Llama3, and as such, might have been pre-exposed to the model during training. We detail our reasons as to why data leakage is not a concern for the conclusions that we draw from this study:
  • There are datasets in our study that have not been leaked. For example, the MMLU test set is used as an official test benchmark by LLama3[26], and WildJailbreak was released to the public only after the initial release of Llama3 models [45]. The results are strong on such datasets as well, showing that data leakage is not behind the performance of the probes.
  • Several tasks require the probes to identify behavior that fundamentally does not exist in text form on the internet. The abstention failure, format following (bullets and JSON) and confidence estimation tasks require the probes to determine how the model will behave, regardless of whether the model has seen the input text or not. For example, whether or not the model was pretrained on MSMarco questions is immaterial for predicting whether or not it will output answers in a specific JSON format.
@@ Appendix B Extended Analysis
B.1 Training Datapoint Ablation
Figure 9: Within 500 training datapoints, probes often approach the consistency and coverage ($\alpha=0.9$) they will attain when trained on the entire dataset.
As a general trend (Figure 9), 500 training datapoints are usually sufficient for probes to achieve the maximum consistency and coverage that they will attain when trained on the entire dataset.
B.2 Layer Ablation
@@
B.3 Confidence Threshold ablation
Figure 11: Increasing the confidence threshold $\alpha$ has no effect until a point, after which conformal consistency increases while the coverage tends to zero.
As a general trend (Figure 11), increasing the confidence threshold $\alpha$ has no impact until a point, after which it leads to a steady decline in coverage and a steady increase in estimation consistency.
@@ C.1.5 Toxicity Detection:
We use two datasets provided by the same organization, JigsawToxicity and JigsawUnintendedBiasToxicity [1]. Both datasets have scores for toxicity from a chatroom setting.
@@ C.2.1 Text Classification Tasks:
MCQ: For the MCQ tasks, we select two options by randomly sampling an incorrect answer from the provided options along with the correct option. This is because the MCQ datasets vary in terms of number of MCQ options, and for the out-of-distribution experiment we require all the datasets to have the same number of answer classes (note that the topic identification task below is multiclass, showing that our method is not limited to binary classification). The CoT model is asked to identify the correct answer option.
@@ C.2.2 Other Tasks:
LM Abstention: For the unanswerable question datasets of SelfAware and KnownUnknown, the model is first instructed to give a reasoning chain about the question, and answer only if the question is answerable (abstaining otherwise). With WildJailbreak, the input need not be a question, but any request. Hence, we instruct the model to comply with the request if it is not malicious, and abstain otherwise. We keep only the instances where the LM responds or complies, and train probes to detect whether or not the LM should have abstained (i.e. unanswerable question or malicious request).
@@ C.3 Prompts and Inference:
Each task and dataset has a different set of prompts. We have provided a link to the code, and all prompts can be seen in the file data.py
All LM inference uses greedy decoding and is hence deterministic.
@@ C.4 Hardware:
All of our experiments were run on a compute cluster with 8 NVIDIA A40 GPUs (approx 46068 MiB of memory) on CUDA version 12.6. The CPU on the cluster is an AMD EPYC 7502 32-Core Processor. Most experiments could be conducted with less than 16GB of GPU RAM.
Generating the hidden states took around 2 hours for every 10,000 points, while training a hidden state probe takes fewer than 30 seconds.